# Inverse Transition Learning: Learning Dynamics from Demonstrations

### v1

@@ Introduction
In traditional planning scenarios, the rewards $R$ and transition dynamics $T^{*}$ of the environment are known, and the goal is to compute the optimal policy $\pi^{*}$ that maximizes long-term returns. However, in many real-world situations, the transition dynamics $T^{*}$ are unknown. Model-Based Reinforcement Learning (MBRL) addresses this by first learning $T^{*}$ and then performing planning, which improves data efficiency and enables counterfactual reasoning (Sutton and Barto 2018; Ghavamzadeh et al. 2015; Kidambi et al. 2020; Yu et al. 2021; Lee, Lee, and Kim 2021; Poupart et al. 2006; Ha and Schmidhuber 2018; Oh et al. 2015; Buesing et al. 2018).
@@
This paper focuses on learning the true dynamics $T^{*}$ in offline settings using batch data generated by a near-optimal expert. This is a common setting in fields like healthcare and education, where one can presume that the behavior policy is imperfect but generally reasonable. Learning $T^{*}$ from observational data is challenging due to low coverage of the state-action space. Without the ability to interact with the environment, it is crucial to utilize the limited data effectively.
We leverage the knowledge that the trajectories are near-optimal to better estimate $T^{*}$. Consider the simple 2-state MDP shown on the next page. The blue path shows the observed optimal behavior leading to the goal state $s_{\text{goal}}$ from $s_{1}$ after action $a_{1}$. The dashed lines represent the alternative hypothetical paths following an unobserved action $a_{2}$. The fact that the expert chose action $a_{1}$ implies there must be a higher likelihood of reaching $s_{\text{goal}}$ via $a_{1}$ than $a_{2}$ (i.e., $T^{*}(s_{\text{goal}}|s_{1},a_{1})>T^{*}(s_{\text{goal}}|s_{1},a_{2})$.
$s_{1}$$s_{\text{goal}}$$a_{1}$$a_{2}$$a_{2}$bserved expert: $a_{1}$nobserved: $a_{2}$
@@
A popular definition of near-optimality models an expert’s actions as being proportional to the action’s values, as presented in Maximum Causal Entropy Inverse (MCE) RL (Ziebart, Bagnell, and Dey 2010).
@@
In clinical settings, clinicians often prefer to administer treatments from drug families that have similar, reasonably good effects, without assigning a specific ranking. Conversely, they avoid poorly performing treatments, also without assigning a ranking. Instead of selecting a single best action, which may be influenced by data collection or other artifacts, it is often better to focus on sets of actions that are nearly equivalent in performance (Tang et al. 2020) and identify which poor behaviors to avoid (Fatemi et al. 2021; Tang et al. 2022; Rebello et al. 2023).
@@
The most closely related works to ours (Herman et al. 2016; Reddy, Dragan, and Levine 2018) rely on the Maximum Causal Entropy (MCE) modeling approach. While modeling near-optimal expert behavior this way is reasonable, our modeling approach shows better performance in clinical settings. In addition, solving for $T^{*}$ within the MCE framework requires a gradient-based procedure within an alternating optimization process, which is computationally intensive and prone to getting stuck in local optima. Additionally, the constraints used in these methods are soft, meaning their learned dynamics do not have any guarantees.
@@
To address these limitations, we introduce a novel approach called Inverse Transition Learning (ITL), based on hard constraints on the true dynamics $T^{*}$ and the distinction between the two groups of actions. Our constraints ensure that the values of executed actions exceed the values of actions not taken. If multiple actions are taken in the same state, they are constrained to be close in value. These thresholds are governed explicitly by a transparent and human-understandable parameter, rather than implicitly through the entropy of an action-value distribution. They can also be solved for deterministically via a quadratic program (e.g., using CVXPY (Diamond and Boyd 2016)), avoiding the drawbacks of gradient-based optimization. This results in an optimization procedure that provides guarantees for the estimated dynamics and converges significantly faster.
@@
The fact that the trajectories are near-optimal provides some information about what transition dynamics $T^{*}$ are feasible, but limited coverage in the batch setting means that some uncertainty will still remain. Thus, we extend our approach to the Bayesian model-based setting (Dearden, Friedman, and Andre 2013). We develop an efficient approach to posterior estimation that includes our constraints. Our approach narrows the gap to the true dynamics $T^{*}$ compared to baseline Bayesian MBRL methods. Additionally, given a new reward function, we demonstrate how keeping a posterior over the true dynamics $T^{*}$ can be used to predict when the transfer will be successful. Our main contributions are:
@@
Related Works
Model-Based Reinforcement Learning. We explore the utilization of forward models, which predict the subsequent state $s^{\prime}$ from the current state $s$ and action $a$ ($(s,a)\rightarrow s^{\prime}$), as defined in (Moerland et al. 2023). These models are central for understanding action-induced state transitions. In contrast, backward models identify potential antecedents to states ($s^{\prime}\rightarrow(s,a)$) and are used for backward planning (Moore and Atkeson 1993). Similarly, inverse models compute actions required to transition between states $(s,s^{\prime}\rightarrow a)$, beneficial in RRT planning (LaValle 1998). Additionally, non-parametric methods like replay buffers enable precise estimations (Lin 1992; Vanseijen and Sutton 2015; Van Hasselt, Hessel, and Aslanides 2019), and approximation methods, such as Gaussian processes, offer alternatives (Wang, Hertzmann, and Fleet 2005; Deisenroth and Rasmussen 2011). In discrete MDP settings like ours, tabular maximum likelihood estimation models, noted as $T^{MLE}$, represent the state of the art (Sutton 1991).
@@
Bayesian MBRL integrates uncertainty into model learning, as detailed in the foundational works (Ghavamzadeh et al. 2015; Ross and Pineau 2008; Dearden, Friedman, and Andre 2013). Unlike previous approaches, such as (Poupart and Vlassis 2008) which address partially observable settings in discrete factored domains by updating beliefs based on new data, our method applies Bayesian MBRL in an offline setting. We infer posterior distributions over transition dynamics solely from expert knowledge. This approach utilizes prior knowledge and batch data to develop models that establish reliable policies (Guo, Yunfeng, and Geng 2022). In contrast, studies like (Zhang et al. 2020a, b) focus on learning invariant representations for control without explicitly modeling transition dynamics, which is a key aspect of our research.
@@
Learning from Demonstrations. Imitation Learning (IL) involves learning policies directly from demonstrations by near-optimal experts using methods like training a policy at each timestep (Stéphane, Gordon Geoffrey, and Andrew 2010), DAgger (an iterative algorithm integrating expert feedback) (Ross, Gordon, and Bagnell 2011), and Approximate Policy Iteration with Demonstration (APID), which uses expert advice to impose linear constraints during policy optimization (Kim et al. 2013). These techniques primarily focus on behavioral cloning and policy learning. In contrast, Inverse Reinforcement Learning (IRL) seeks to learn the reward function $R$ from the dynamics $T$ and expert demonstrations, facilitating succinct task descriptions and transferability (Ng, Russell et al. 2000). Methods like Maximum Margin and Maximum Entropy IRL focus on deriving $R$ to mimic expert policies robustly (Abbeel and Ng 2004; Ziebart et al. 2008; Scobee and Sastry 2019). Addressing $R$’s non-identifiability, Bayesian IRL suggests learning a distribution over $R$, enhancing model adaptability (Ramachandran and Amir 2007). Distinct from these approaches, our work Inverse Transition Learning emphasizes learning the true dynamics $T^{*}$, a novel framework constructed to integrates expert demonstrations to refine the estimation of $T^{*}$. This method aims to understand and model the actual transitions within the environment, contrasting sharply with baseline methods that typically rely on maximum likelihood for estimation.
Preliminaries
Markov Decision Processes (MDPs). An MDP $\mathcal{M}$ can be represented as a tuple =$\mathcal{M}=\{\mathcal{S},\mathcal{A},T^{*},\gamma,R\}$ , where $\mathcal{S}$ is the state space , $\mathcal{A}$ the action space, $T^{*}$ is the true dynamics of the environment, $\gamma$ is the discount factor and $R$ is a bounded reward function. In planning, the goal is to find the best policy $\pi^{*}$ corresponding to an MDP $\mathcal{M}$. In this paper, we focus on tabular MDPs (discrete state and action spaces).
Bellman Equations. The values functions of a policy $\pi$ are given below:
$V^{\pi}(s)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\mid s_{0}=s,\pi\right]$ (1)
$Q^{\pi}(s,a)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\mid s_{0}=s,a_{0}=a,\pi\right]$ (2)
@@
Notation. Let $V^{\pi}\in\mathbb{R}^{|S|}$ denote the vector of values $V^{\pi}(s)$. We use shorthand $R_{a}\in\mathbb{R}^{|S|}$, $Q_{a}\in\mathbb{R}^{|S|}$, and $T_{a}\in\mathbb{R}^{|S|\times|S|}$ to represent the vectors $R(\cdot,a)$, $Q(\cdot,a)$ and the matrix $T(\cdot\mid\cdot,a)$. We also use shorthand $R_{\pi},Q_{\pi}$, and $T_{\pi}$ to represent the vectors $\mathbb{E}_{a\sim\pi}[R_{a}]$. $\mathbb{E}_{a\sim\pi}[Q_{a}]$ and the matrix $\mathbb{E}_{a\sim\pi}[T_{a}]$. For dynamics $T$, $\pi^{*}(T)$ denotes its corresponding optimal policy. In tabular settings, the value functions $V^{\pi}$ and $Q^{\pi}_{a}$ can be calculated directly through a closed-form solution:
@@
Methods for Estimating the Transition Dynamics $T^{*}$. In tabular settings, a common method to estimate the true transition dynamics $T^{*}$ is the Maximum Likelihood Estimate (MLE), denoted as $T^{MLE}$ (e.g., (Barto, Bradtke, and Singh 1995; Kim and Oh 2023; Ornik and Topcu 2021)). The number of times the tuple (s, a, s’) is observed is denoted by $N_{s,a,s^{\prime}}$. We apply Laplace smoothing, represented by $\delta$, to address issues such as zero occurrences in the count data due to low coverage in our batch data $\mathcal{D}$. The $T^{MLE}$ is defined as:
@@
In offline tabular settings, (Herman et al. 2016) leverage near-optimal batch data to infer dynamics by reusing the MCE IRL framework (Ziebart, Bagnell, and Dey 2010). We include this method in our baseline and refer to it as MCE (or $T^{MCE}$) moving forward. Their method consist of an interative procedure between taking one gradient step of the dynamics parameters $\theta$ with respect to the loss in equation 6 and then performing soft-Value Iteration, where Q is the soft Q function.
$L_{\theta}^{MCE}(D)=\sum_{(s,a,s^{\prime})\in D}$ $\left[\log\left(\frac{\exp\left(Q_{\theta}(s,a)\right)}{\sum_{a^{\prime}\in\mathcal{A}}\exp\left(Q_{\theta}(s,a^{\prime})\right)}\right)\right.$ (6)
$\left.+\log T_{\theta}\left(s^{\prime}\mid s,a\right)\right].$
To model the uncertainty of transitions $T^{*}$ probabilistically, (Ghavamzadeh et al. 2015) utilizes a Multinomial likelihood function in conjunction with a Dirichlet prior. This combination yields the posterior distribution $P(T^{*}|\mathcal{D})$, which estimates the transition probabilities based on the observed data $\mathcal{D}$.
$P(T^{*}(\cdot\mid s,a)\mid\mathcal{D})=\text{Dir}(\mathbf{N}_{s,a}+\mathbf{\delta}\mid s,a)\\ \propto\text{Multinomial}(\mathbf{N}_{s,a}\mid s,a)\cdot\text{Dir}(\mathbf{\delta}\mid s,a)$ (7)
Note that $T^{MLE}$ represents the mean of the distribution $P(T^{*}|\mathcal{D})$. However, this posterior distribution is not as tight as it could be because it does not account for the near-optimality of expert trajectories. To refine this, we utilize expert signals to develop a tighter posterior $P(T^{*}|\mathcal{D})$, enhancing the estimation of the transition dynamics. To the best of our knowledge, we are the first to infer a distribution over the dynamics using expert demonstrations. Furthermore, we propose a method distinct from (Herman et al. 2016) to derive a point estimate of $T^{*}$. This involves formulating constraints based on expert behaviors, offering a more accurate approach to modeling the true dynamics.
Estimating Transition Dynamics with $\epsilon$-optimal Expert
This section introduces key definitions to relate the expert’s optimality to the true dynamics $T^{*}$. We describe the problem setting as follows:
Definition 1.
($\epsilon$-ball) For any state $s$ and transition dynamics $T$, an action $a$ is in the $\epsilon$-ball $\epsilon(s;T)$ if it is $\epsilon$-close to the optimal action according to the optimal Q-function $Q^{*}(.,.;T)$ (We refer to actions as valid if they are within the $\epsilon$-ball for state $s$ and dynamics $T$, and as invalid if they are outside this $\epsilon$-ball):

### v2

# Bayesian Inverse Transition Learning: Learning Dynamics From Near-Optimal Trajectories

@@ 1 INTRODUCTION
In traditional planning scenarios, the rewards $R$ and transition dynamics $T^{*}$ of the environment are known, and the goal is to compute the optimal policy $\pi^{*}$ that maximizes long-term returns. However, in many real-world situations, the transition dynamics $T^{*}$ are unknown. Model-Based Reinforcement Learning (MBRL) addresses this by first learning $T^{*}$ and then performing planning, which improves data efficiency and enables counterfactual reasoning (Sutton and Barto,, 2018; Ghavamzadeh et al.,, 2015; Kidambi et al.,, 2020; Yu et al.,, 2021; Lee et al.,, 2021; Poupart et al.,, 2006; Ha and Schmidhuber,, 2018; Oh et al.,, 2015; Buesing et al.,, 2018).
@@
This paper focuses on learning the true dynamics $T^{*}$ in offline settings using batch data generated by a near-optimal expert. This is a common setting in fields like healthcare and education, where one can presume that the behavior policy is imperfect but generally reasonable. Learning $T^{*}$ from observational data is challenging due to low coverage of the state-action space. Without the ability to interact with the environment, it is crucial to utilize the limited data effectively. We leverage the knowledge that the trajectories are near-optimal to better estimate $T^{*}$. Consider the simple 2-state MDP shown below. The blue path shows the observed optimal behavior leading to the goal state $s_{\text{goal}}$ from $s_{1}$ after action $a_{1}$. The dashed lines represent the alternative hypothetical paths following an unobserved action $a_{2}$. The fact that the expert chose action $a_{1}$ implies there must be a higher likelihood of reaching $s_{\text{goal}}$ via $a_{1}$ than $a_{2}$ (i.e., $T^{*}(s_{\text{goal}}|s_{1},a_{1})>T^{*}(s_{\text{goal}}|s_{1},a_{2})$.
$s_{1}$$s_{\text{goal}}$$a_{1}$$a_{2}$$a_{2}$Observed expert: $a_{1}$Unobserved: $a_{2}$
@@
A popular definition of near-optimality models an expert’s actions as being proportional to the action’s values, as presented in Maximum Causal Entropy (MCE) Inverse RL (Ziebart et al.,, 2010).
@@
In clinical settings, clinicians often prefer to administer treatments from drug families that have similar, reasonably good effects, without assigning a ranking. Conversely, they avoid poorly performing treatments, also without assigning a ranking. Instead of selecting a single best action, which may be influenced by data collection or artifacts, it is often better to focus on sets of actions that are nearly equivalent in performance (Tang et al.,, 2020) and identify which poor behaviors to avoid (Fatemi et al.,, 2021; Tang et al.,, 2022).
@@
The most closely related works to ours (Herman et al.,, 2016; Reddy et al.,, 2018) rely on the Maximum Causal Entropy (MCE) approach. While modeling near-optimal expert behavior this way is reasonable, our approach shows better performance in clinical settings. Moreover, solving for $T^{*}$ in the MCE framework requires a gradient-based alternating optimization, which is computationally intensive and prone to local optima. Finally, the constraints in these methods are soft, so their learned dynamics lack guarantees.
@@
To address these limitations, we introduce a novel approach called Inverse Transition Learning (ITL), based on hard constraints on the true dynamics $T^{*}$ and the distinction between two groups of actions. Our constraints ensure that executed actions have higher value than actions not taken. If multiple actions are taken in the same state, they are constrained to be close in value. These thresholds are governed explicitly by a transparent, human-understandable parameter, rather than implicitly through the entropy of an action-value distribution. They can also be solved deterministically via a quadratic program (e.g., using CVXPY (Diamond and Boyd,, 2016)), avoiding the drawbacks of gradient-based optimization. This yields an optimization procedure that provides guarantees for the estimated dynamics and converges significantly faster.
@@
The fact that the trajectories are near-optimal provides some information about what transition dynamics $T^{*}$ are feasible, but limited coverage in the batch setting means that some uncertainty will still remain. Thus, we extend our approach to the Bayesian model-based setting (Dearden et al.,, 2013). We develop an efficient approach to posterior estimation that includes our constraints. Our approach narrows the gap to the true dynamics $T^{*}$ compared to baseline Bayesian MBRL methods. Additionally, given a new reward function, we demonstrate how keeping a posterior over the true dynamics $T^{*}$ can be used to predict when the transfer will be successful. Our main contributions are:
@@
2 RELATED WORKS
Model-Based Reinforcement Learning. We explore the utilization of forward models, which predict the subsequent state $s^{\prime}$ from the current state $s$ and action $a$ ($(s,a)\rightarrow s^{\prime}$), as defined in (Moerland et al.,, 2023). These models are central for understanding action-induced state transitions. In contrast, backward models identify potential antecedents to states ($s^{\prime}\rightarrow(s,a)$) and are used for backward planning (Moore and Atkeson,, 1993). Similarly, inverse models compute actions required to transition between states $(s,s^{\prime}\rightarrow a)$, beneficial in RRT planning (LaValle,, 1998). Additionally, non-parametric methods like replay buffers enable precise estimations (Lin,, 1992; Vanseijen and Sutton,, 2015; Van Hasselt et al.,, 2019), and approximation methods, such as Gaussian processes, offer alternatives (Wang et al.,, 2005; Deisenroth and Rasmussen,, 2011). In discrete MDPs like ours, tabular maximum likelihood estimation models, noted as $T^{MLE}$, represent the state of the art (Sutton,, 1991).
@@ 2 RELATED WORKS
Bayesian MBRL integrates uncertainty into model learning, as detailed in foundational works (Ghavamzadeh et al.,, 2015; Ross and Pineau,, 2008; Dearden et al.,, 2013). Unlike earlier approaches, such as (Poupart and Vlassis,, 2008), which address partially observable settings in discrete factored domains by updating beliefs with new data, our method applies Bayesian MBRL in an offline setting. We infer posterior distributions over transition dynamics solely from expert knowledge. This approach combines prior knowledge with batch data to develop models that yield reliable policies (Guo et al.,, 2022). In contrast, studies like (Zhang et al., 2020a, ; Zhang et al., 2020b, ) focus on learning invariant representations for control without explicitly modeling transition dynamics.
@@
Learning from Demonstrations. Imitation Learning (IL) learns policies directly from expert demonstrations via per-timestep training (Stéphane et al.,, 2010), iterative expert feedback (DAgger) (Ross et al.,, 2011), or imposing linear constraints during policy optimization (APID) (Kim et al.,, 2013). In contrast, Inverse Reinforcement Learning (IRL) infers the reward $R$ from dynamics $T$ and demonstrations to facilitate transferability (Ng et al.,, 2000), using methods like Maximum Margin and Maximum Entropy to robustly mimic experts (Abbeel and Ng,, 2004; Ziebart et al.,, 2008; Scobee and Sastry,, 2019). To address $R$’s non-identifiability, Bayesian IRL learns a distribution over $R$ (Ramachandran and Amir,, 2007). Distinct from these policy and reward-focused approaches, our Inverse Transition Learning framework utilizes expert demonstrations to directly refine the estimation of the true dynamics $T^{*}$, contrasting with standard baselines that rely solely on maximum likelihood.
3 PRELIMINARIES
Markov Decision Processes (MDPs). An MDP $\mathcal{M}$ can be represented as a tuple $\mathcal{M}=\{\mathcal{S},\mathcal{A},T^{*},\gamma,R\}$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $T^{*}$ is the true dynamics of the environment, $\gamma$ is the discount factor, and $R$ is a bounded reward function. In planning, the goal is to find the best policy $\pi^{*}$ corresponding to an MDP $\mathcal{M}$. In this paper, we focus on tabular MDPs (discrete state and action spaces). Discretization of continuous clinical variables can be achieved using domain knowledge and is a common and effective practice in healthcare RL applications. This approach aligns naturally with clinical decision-making, where physicians often categorize patient states and treatment options into discrete, clinically meaningful groups. For example, (Choudhary et al.,, 2024) use the same healthcare dataset as us (MIMIC (Johnson et al.,, 2020)) to create an RL simulator by discretizing both state and action spaces. Such discretization not only reflects clinical practice but also makes the learned dynamics more interpretable in low-data settings.
Value Functions. For a policy $\pi$, the state–value function $V^{\pi}(s)$ is the expected discounted return starting from state $s$, and the action–value function $Q^{\pi}(s,a)$ is the expected discounted return starting from state $s$ and taking action $a$.
@@ 3 PRELIMINARIES
Notation. Let $V^{\pi}$ denote the vector of values $V^{\pi}(s)$. We use shorthand $R_{a}$ , $Q_{a}$, and $T_{a}$ to represent the vectors $R(\cdot,a)$, $Q(\cdot,a)$ and the matrix $T(\cdot\mid\cdot,a)$. We also use shorthand $R_{\pi},Q_{\pi}$, and $T_{\pi}$ to represent the vectors $\mathbb{E}_{a\sim\pi}[R_{a}]$. $\mathbb{E}_{a\sim\pi}[Q_{a}]$ and the matrix $\mathbb{E}_{a\sim\pi}[T_{a}]$. For dynamics $T$, we let $\pi^{*}(T)$ and $Q^{*}(T)$ denote the corresponding optimal policy and Q values function respectively with respect to dynamics $T$. In tabular settings, the value functions $V^{\pi}$ and $Q^{\pi}_{a}$ can be calculated directly through a closed-form solution:
@@
Methods for Estimating the Transition Dynamics $T^{*}$. In tabular settings, a common method to estimate the true transition dynamics $T^{*}$ is the Maximum Likelihood Estimate (MLE), denoted as $T^{MLE}$ (e.g., (Barto et al.,, 1995; Kim and Oh,, 2023; Ornik and Topcu,, 2021)). The number of times the tuple (s, a, s’) is observed is denoted by $N_{s,a,s^{\prime}}$. We apply Laplace smoothing, represented by $\delta$, to address issues such as zero occurrences in the count data due to low coverage in our batch data $\mathcal{D}$. The $T^{MLE}$ is defined as:
@@
In offline tabular settings, (Herman et al.,, 2016) leverage near-optimal batch data to infer dynamics by reusing the MCE IRL framework (Ziebart et al.,, 2010). We include this method in our baseline and refer to it as MCE (or $T^{MCE}$) moving forward. Their method consist of an interative procedure between taking one gradient step of the dynamics parameters $\theta$ with respect to the loss in equation 4 and then performing soft-Value Iteration, where Q is the soft Q function.
$\sum_{(s,a,s^{\prime})\in\mathcal{D}}\left[\log\frac{\exp(Q_{\theta}(s,a))}{\sum_{a^{\prime}}\exp(Q_{\theta}(s,a^{\prime}))}+\log T_{\theta}(s^{\prime}\mid s,a)\right].$ (4)
To model the uncertainty of transitions $T^{*}$ probabilistically, (Ghavamzadeh et al.,, 2015) uses a Multinomial likelihood in conjunction with a Dirichlet prior. This combination yields the posterior $P(T^{*}|\mathcal{D})$, which estimates the transition probabilities given data $\mathcal{D}$.
$P(T^{*}(\cdot\mid s,a)\mid\mathcal{D})$ $=\text{Dir}(\mathbf{N}_{s,a}+\mathbf{\delta}\mid s,a)$ (5)
$\propto\text{Multinomial}(\mathbf{N}_{s,a}\mid s,a)\cdot\text{Dir}(\mathbf{\delta}\mid s,a)$
Note that $T^{MLE}$ represents the mean of the distribution $P(T^{*}|\mathcal{D})$. However, this posterior distribution is not as tight as it could be because it does not account for the near-optimality of expert trajectories. To refine this, we utilize expert signals to develop a tighter posterior $P(T^{*}|\mathcal{D})$, enhancing the estimation of the transition dynamics. To the best of our knowledge, we are the first to infer a distribution over the dynamics using expert demonstrations. Furthermore, we propose a method distinct from (Herman et al.,, 2016) to derive a point estimate of $T^{*}$. This involves formulating constraints based on expert behaviors, offering a more accurate approach to modeling the true dynamics.
4 ESTIMATING TRANSITION DYNAMICS WITH $\epsilon$-OPTIMAL EXPERT
This section introduces key definitions to relate the expert’s optimality to the true dynamics $T^{*}$.
Definition 1 ($\epsilon$-ball).
For any state $s$ and transition dynamics $T$, an action $a$ is in the $\epsilon$-ball $\epsilon(s;T)$ if it is $\epsilon$-close to the optimal action according to its optimal Q-function $Q^{*}(.,.;T)$ with respect to dynamics $T$: